# “Rich-Get-Richer”? Analyzing Content Creator Earnings Across Large Social Media Platforms

### v1

@@
“Rich-Get-Richer”? Analyzing Content Creator Earnings Across Large Social Media Platforms
@@
This paper examines whether monthly content creator earnings follow a power law distribution, driven by compounding ‘rich-get-richer’ dynamics [6]. Patreon creator earnings data for 2018, 2021, and 2024 for Instagram, Twitch, YouTube, Twitter, Facebook, and Patreon exhibit a power law exponent around $\alpha=2$. This suggests that algorithmic systems generate unequalizing returns closer to highly concentrated capital income and wealth, rather than labor income. Platforms governed by powerful and compounding recommendation systems, such as Instagram and YouTube, exhibit both a stronger power law relation (lower $\alpha$) and lower mean, median, and interquartile earnings, indicating algorithms that disproportionately favor top earners at the expense of a ‘middle class’ of creators. In contrast, Twitter and Patreon have a more moderate $\alpha$, with less earnings inequality and higher middle class earnings. Policies which incentivize the algorithmic promotion of longer-tail content (to explore more and exploit less) may help creator ecosystems become more equitable and sustainable.
@@ 1  Introduction
Despite their significance, digital labor markets and their associated earnings distributions remain largely opaque and unregulated [49, 33]. This paper aims to fill this gap by examining how algorithms shape earnings distributions in digital platforms and the welfare implications of these patterns. The declining quality of platforms’ algorithmic allocations has been an important reflection of their increasing market power [59, 50, 19]. But the relationship between earnings allocations within platforms – driven by algorithmic allocations of user attention – and market structure is not straightforward (Section 4).
@@
One useful way to understand and characterize earnings online is the power law distribution, $P(X>x)\sim\frac{1}{x^{\alpha}}$, which can provide insights into the scaling of the system, the extent of earnings inequality, and the possible mechanisms behind it. Power laws describe distributions where small occurrences are extremely common, while large occurrences are extremely rare but have a significant impact. A lower power law exponent generally implies a higher degree of inequality, since it means a greater probability of finding very high (extreme) tail incomes, characteristic of long-tail distributions such as wealth [51, 3]. Long-tail distributions famously occur online too, such as with internet traffic, where a few websites get the majority of visits [1, 37].
Several factors have been found to reinforce the unequal dynamics which network effects create online, chief among them algorithmic selection and recommendation of content, which frequently amplifies pre-existing popularity [66, 36], creating a ‘winner takes all’ or viral effect [57, 56]. This compounding of initial advantage, also called a ‘superstar effect’ in economics [28, 4], can be characterized more generally by a ‘preferential attachment’ model [6]. The hypothesis we estimate is that the same ‘rich-get-richer’ dynamics that drive attention in the preferential attachment model [6] also govern content creator earnings. As creators gain visibility, they attract further algorithmic attention, leading to a large concentration of income among a small group of highly popular creators. Earnings data, therefore, provide a concrete and measurable outcome of these dynamics. A similar dynamic of persistence has been shown to govern firm dynamics in the economy [31, 30, 10], however, online, network and algorithmic effects can supercharge this [21].11endnote: 1The low marginal cost to distributing and selling content online adds to the power law, which arises from the feedback introduced by correlated decisions across a population [21].
The key questions we investigate are whether the distribution of Patreon content creator earnings across major social media platforms follows a power law pattern, and how the estimated power law exponent ($\alpha$) reflects the health and equity of the platform ecosystem.
We answer this question by exploiting a novel feature in a large dataset (n = 104,719) of Patreon monthly earnings data for the years 2018, 2021, and 2024, which allows us to identify creators who use only a single social media site (or none). Patreon is a monetization platform for content creators, allowing them to sell additional or primary content to subscribers (paid and unpaid). Earnings here are a direct reflection of popularity, as monetization mechanisms – such as ad revenue sharing, sponsorships, and promotions – are closely linked to a creator’s reach and influence.
We find that the monthly earnings distribution of Patreon content creators across platforms fits a power law distribution with an exponent around $\alpha=2.0$ (lower exponent is more concentrated), highlighting an influential tail of a few high earning content creators. A powerful ‘rich-get-richer’ dynamic, represented by a preferential attachment model [6], is an excellent fit for explaining how this power law arises (as the limiting distribution) on YouTube ($\alpha=1.8$), Instagram ($\alpha=1.84$), Twitch ($\alpha=1.93$), and Facebook ($\alpha=1.94$).22endnote: 2For $\alpha<2$ the mean of the distribution converges, which means the average earnings across creators is finite. However, the variance diverges, indicating that earnings can vary wildly, with no typical range for large creators. Practically, this means that the mechanism(s) generating content creator earnings online [59] is closer to the multiplicative and compounding effects of capital and wealth ($\alpha=1.5$ typically) [28, 42].33endnote: 3For debate see: [12].
Secondly, we find that the ‘rich-get-richer’ dynamic, characteristic of platforms with stronger power laws (lower $\alpha$), may harm the platform’s ‘middle class’ of content creators, as they also land up experiencing lower mean and median earnings, with a less expansive interquartile earnings range (Q25-Q75). In contrast, platforms with weaker power law relations – namely Patreon and especially Twitter – exhibit a somewhat healthier ecosystem that does support a middle-class of creator earnings. This may be driven by the potentially compounding nature and centralized integration of algorithmic content recommendation systems on platforms such as Instagram and YouTube [18, 17], which may not sufficiently promote lesser-known creators [2]. Patreon’s more moderate $\alpha$ value may be due to its direct creator-audience relationships. Twitter’s weaker power law reflects more opportunities for content creators, especially ‘NSFW’ (‘not safe for work’) creators earning below $2,114$ a month. But for those earning above this, Twitter’s earning opportunities are in fact more concentrated that on other platforms.
Thirdly, we find a growing convergence (and worsening) over time in YouTube, Twitch, Instagram, and Facebook power laws, perhaps as algorithmic influence grows (relative to social graph mechanisms), $\alpha$ values decline, implying growing concentration in earnings among a few. Instagram shows the most extreme change, with $\alpha$ worsening from around 2.1 in 2021 or above to around 1.84 in 2024. Patreon content creators on Twitter appear not to have suffered from Elon Musk’s takeover of the platform in October 2022, with their earnings distribution becoming less unequal (higher $\alpha$) relative to 2021. The main impact in the data is that far more NSFW content creators left Instagram and joined Twitter since 2021. But, overall, there are no countervailing forces to offset pure ‘rich-get-richer’ (preferential attachment) dynamics on platforms. Potential corrective measures (Section 4), such as mandated use and disclosure of recommendation systems by platforms that better incentivize the creation and reward of longer-tail content [52, 35, 65], may be the most effective means of fostering a more equitable ecosystem of content creators online. And to the extent that more unequal earnings distributions reflects platform market power, rather than more inter-platform competition, antitrust policy has a role to play too (Section 4).
Notably, unlike income distribution models where the power law typically governs only the extreme right tail, with the majority of the distribution better explained by exponential or gamma-type distributions [61], our results indicate that social media earnings data as a whole are largely dominated by a power law, as evidenced by the small estimated minimum dollar values at which power law dynamics emerge. Despite a few previous studies employing Patreon earning data [55, 24, 38], we provide the first systematic analysis of the nature of the distribution – and in turn the likely economic process or algorithmic mechanisms generating this data. Moreover, we provide some of the only evidence on the distribution of earnings of content creators [40, 53].
A key limitation is the use of Patreon content creators earnings as a proxy for the health of platform content creators as a whole, given that the underlying monetization mechanisms do differ somewhat. This means that we cannot rule Patreon specific drivers for our findings (such as users having to click through to the Patreon creator’s site) or a selection bias at play. However, the weaker power law dynamic operating on Patreon creators active on Twitter, and those without alternative active social media platforms (content creators reliant only on Patreon), greatly reduces the veracity of this argument. In the following Section 2 we lay out our data and power law method, along with the underlying ‘rich-get-richer’ dynamic underpinning it – discussed further in Appendices. This is followed by our results (Section 3) and a concluding section with a brief discussion on market power and recommender system types, including how to incentivize greater incorporation of non-viral (longer-tail) content [2] in recommender systems (Section 4).
@@ 2  Data and Method
This section motivates our use of a power law to estimate earnings, and our interpretation of it as reflecting a ‘rich-get-richer’ dynamic, drawing on the preferential attachment model outlined in [6] (‘BA’ model). This ‘BA’ model predicts a power law distribution of attention with an exponent $\alpha=2$ in cases of pure preferential attachment. Before discussing the model we detail our data.
@@ 2.1  Data and Descriptive Statistics
We use official Patreon monthly earnings data for content creators purchased directly from the Patreon platform, for the month of March for years 2018, 2021, and 2024. Significant earnings data is missing, especially for larger earners (since those with more paid subscribers tend to disclose their earnings the least). Not imputing missing earnings, therefore, would introduce significant downward bias on the right tail of the distribution. We impute missing earnings linearly using a basic specification, based on earnings being a function of paid members (subscribers), all members, the type of content created, if they are an NSFW content creator, and the year of the data. This is unlikely to the sole type of earnings for these content creators, since they can also make money directly from their affiliation with the social media platform, e.g. from YouTube ads and marketing, unless their only affiliation is Patreon. We restrict the dataset to include only content creators with monthly earnings exceeding \$10 USD.
We segment the data by the social media platform(s) the Patreon content creators use to engage their audiences. We identify creators who only use a single social media platform and remove from our dataset all earners with more than one social media affiliation. The dataset spans Patreon creators across major platforms, including YouTube, Instagram, Twitter, Twitch, Facebook, and Patreon only. Table 1 provides summary statistics for earnings by platform, detailing the number of observations, mean earnings, standard deviation (SD), and key percentiles (minimum, 25th, 50th, 75th, and maximum).
Table 1: Summary statistics of Patreon monthly earnings by social media platform (US\$, except Obs).
@@
Facebook 8,458 149 47 631 10.00 29.6 46.5 99.0 35,261
Instagram 35,879 226 59 1,076 10.00 33.3 59.3 142.0 60,391
Patreon 83,950 278 57 1,667 10.00 30.7 56.6 154.1 211,321
Twitch 1,793 198 46 1,174 10.04 28.8 46.0 104.6 31,436
Twitter 47,564 373 72 2,159 10.00 33.5 71.9 214.8 169,106
Youtube 23,435 250 47 1,490 10.00 28.4 46.7 118.7 87,802
@@
The data highlights significant variation in earnings across platforms, reflecting differences in audience engagement, platform monetization dynamics, and creator ecosystems. Twitter has the highest mean, median, and interquartile earnings, along with Patreon, while Facebook and Twitch have the lowest mean, median, and interquartile earnings respectively. We note later that this proxy for ecosystem health or strength (median, mean, and interquartile earnings), correlates strongly with the estimated power law for each platform. The standard deviation of earnings is highest for Twitter and lowest for Facebook. Facebook is the only platform to have fewer exclusive Patreon content creators in our data in 2024 than it had in 2021 (Table 2 in Appendix).
@@ 2.2  Power Law Model
Much evidence shows that many aspects of society and the economy follow power law distributions [28], in which the vast majority of outcomes are small, yet a few are extraordinarily large and command a substantial share of the total. Online markets reflect similar patterns, often exhibiting extreme disparities in earnings. For example, a small number of top content creators or successful e-commerce sellers may capture a disproportionate share of total revenue [27, 16].
A power law perspective offers a general framework for understanding the “rich-get-richer” dynamics common in digital platforms [21, 9]. As an entity – be it a creator or a seller – gains initial visibility, algorithms frequently amplify this advantage, leading to further growth in both attention and earnings [66, 36]. This process, known as preferential attachment [6], creates a self-reinforcing feedback loop: as certain participants accumulate more engagement, they become even more likely to attract future interactions, thereby concentrating earnings at the top. By examining earnings distributions and identifying whether they adhere to a power law, researchers can better understand the underlying reinforcement mechanisms and the extent to which disparities are entrenched.44endnote: 4Some econophysics models distinguish a two-class structure in income distributions, where the working class income follows an exponential distribution (reflecting additive processes) and the rentier or capitalist class income follows a power law (reflecting multiplicative processes). See [43, 20].
Detecting a power law distribution in earnings data can, therefore, provide valuable insights into the health and fairness of online ecosystems. Highly skewed distributions, with a large “underclass” earning very little, may signal systemic challenges such as algorithmic bias, barriers to entry, or a lack of upward mobility for newcomers [23, 47]. Conversely, distributions less dominated by a small set of top earners suggest a more inclusive and competitive environment, offering broader opportunities for participants to succeed. Understanding these patterns can inform platform governance, guide policymakers, and help researchers propose fairer algorithms and revenue-sharing mechanisms, ultimately leading to more equitable digital economies [22, 33].

### v2

# “Rich-Get-Richer”? Platform Attention and Earnings Inequality using Patreon Earnings Data

@@
“Rich-Get-Richer”? Platform Attention and Earnings Inequality using Patreon Earnings Data
@@
Using monthly Patreon earnings, we quantify how platform attention algorithms shape earnings concentration across creator economies. Patreon is a tool for creators to monetize additional content from loyal subscribers but offers little native distribution, so its earnings proxy well for the attention creators capture on external platforms (Instagram, Twitch, YouTube, Twitter/X, Facebook, and “Patreon-only”). Fitting power-law tails to test for a highly unequal earnings distribution, we have three key findings. First, across years and platforms the earnings tail and distribution exhibits a Pareto exponent around $\alpha\approx 2$, closer to concentrated capital income than to labor income and consistent with a compounding, “rich-get-richer” dynamic [barabasi1999emergence]. Second, when algorithms tilt more attention toward the top, the gains are drawn disproportionately from the creator “middle class”. Third, over time, creator inequality across social media platforms converge toward similarly heavy-tailed (and increasingly concentrated) distributions, plausibly as algorithmic recommendations rises in importance relative to user-filtered content via the social graph. While our Patreon-sourced data represents a small subset of total creator earnings on these platforms, it provides unique insight into the cross-platform algorithmic effects on earnings concentration.
@@ 1  Introduction
Online social media markets are governed by algorithmic allocations of attention [o2024algorithmic, strauss2024Amazon], not prices.11endnote: 1Pure “matchmakers,” such as Uber, often bypass prices by utilizing search, feed, or recommendation algorithms to allocate resources. Unlike classical multi-sided market models [rochet2003platform, rochet2006two], where prices balance supply and demand, online platforms rely predominantly on non-price algorithmic mechanisms that process vast data on content and user preferences to generate matches and allocate attention [cunningham2024we].
These algorithmic allocations have significant welfare implications. They may disproportionately benefit one side of the platform – users, advertisers, the platform itself, or independent creators – or favor certain types of creators over others [rock2024clicks]. As discovery shifts from user-chosen social graphs to algorithmic recommendations [bhargava2022creator], these attention allocation mechanisms largely determine creator earnings distributions. Platforms that recommend the same “viral” content to all users may generate highly unequal earnings if only a few creators are promoted widely. Conversely, algorithms that surface niche or diverse "long tail" content [anderson2012long] may foster more equitable earnings [brennan2020attention].
Unfortunately, there is little detailed evidence of the economic impact of algorithmic attention mechanisms, since platforms do not disclose creator earnings [gillespie2018custodians, oreilly2023regulating, Naceva_2025_CreatorEP].22endnote: 2Social media platforms are not covered by national statistical classifications, such as the U.S. Census Bureau. Patreon is perhaps the only large creator platform to disclose earnings, but this has not been analyzed systematically to test for cross-platform attention and earnings inequality [regner2021crowdfunding, hutson2021youtube, el2022quantifying, anderson2022patreon].
This paper fills this gap by using monthly Patreon earnings to examine attention dynamics generated by third-party social media platform algorithms, revealing whether these algorithms foster unequal ecosystems or support a healthy creator middle class. We exploit a novel feature in a large multi-year Patreon dataset ($n=104,719$ creators across 2018, 2021, and 2024) that identifies creators who use only a single social media platform (or only Patreon). As a monetization tool for selling additional content to loyal subscribers, Patreon offers little native distribution, making its earnings an ideal proxy for attention captured on external platforms (Instagram, Twitch, YouTube, Twitter/X, Facebook).
@@
We characterize earnings using power law distributions, $P(X)\sim\frac{1}{x^{\alpha}}$, which describe systems where small occurrences are extremely common while rare large occurrences have significant impact. A lower power law exponent implies higher inequality, indicating greater probability of extreme tail incomes characteristic of wealth distributions [pareto1896cours, atkinson1970distribution]. Such heavy-tail distributions also occur in internet traffic, where a few websites capture most visits [adamic2002zipf, huberman1998internet].
Our core finding is that monthly Patreon earnings across platforms fit a power law with exponent around $\alpha\approx 2.0$, indicating an extremely heavy-tailed distribution and a strong “rich-get-richer” dynamic consistent with BA-style preferential attachment and related mixed copying models.[barabasi1999emergence, easley2010networks, chung2003generalizations] This places creator earnings closer to concentrated capital income ($\alpha\approx 1.5$) [gabaix2009power, kumar2024power] than to labor income, suggesting multiplicative and compounding effects drive online creator earnings [strauss2024Amazon].33endnote: 3For debate see: [brzezinski2014wealth]. Platform variation reflects differences in algorithmic systems: YouTube ($\alpha=1.8$), Instagram ($\alpha=1.84$), Twitch ($\alpha=1.93$), and Facebook ($\alpha=1.94$).
Second, platforms with stronger power laws (lower $\alpha$), such as YouTube and Instagram, exhibit weaker creator middle classes: lower mean and median earnings with less expansive interquartile ranges (Q25-Q75). The “rich-get-richer” dynamic draws gains disproportionately from mid-tier creators. In contrast, platforms with weaker power laws—Patreon and especially Twitter—support healthier middle-class creator ecosystems, likely reflecting algorithmic differences in content discovery.
Third, we observe growing convergence toward steeper inequality over time across YouTube, Twitch, Instagram, and Facebook, with declining $\alpha$ values implying increasing earnings concentration. Instagram shows the most extreme shift, with $\alpha$ worsening from around 2.1 in 2021 to 1.84 in 2024, plausibly as algorithmic recommendations rise in importance relative to social-graph mechanisms.
How should we interpret these findings given that Patreon earnings represent only one revenue stream? Our data provides a partial view of total creator earnings inequality and so should not be interpreted as representative of total earnings. That being said, the major creator revenue sources – sponsorships, affiliate marketing, ads – likely exhibit similar (or greater) inequality since they depend on the same attention mechanism: more popular creators earn more [imh_creator_earnings_2023, rieder2023making, oreilly2023regulating].44endnote: 4Patreon monetization likely represents a small component overall. One study finds 2.55% of YouTube video descriptions contain crowdfunding links versus 49% for cross-platform links. Our earnings data most closely proxies for algorithmic attention mechanisms on third-party platforms where these creators are active, since Patreon monetization relies almost entirely on external platforms’ algorithms to generate attention – Patreon offers minimal native distribution.55endnote: 5Patreon’s creator search functionality is primitive. Supporting this interpretation is our finding that creators with no other social media accounts (Patreon-only) show the second least concentrated distribution (highest $\alpha$), after Twitter – whose long reliance on social graphs rather than pure recommendation may explain its relatively equitable outcomes.
The declining quality of platforms’ algorithmic allocations reflects their increasing market power [strauss2024Amazon, o2024algorithmic, Doctorow2024] – though the relationship between within-platform earnings allocations (driven by algorithmic attention) and market structure is not straightforward (Section 5). Potential corrective measures (Section 5), such as mandating recommendation systems that better incentivize long-tail content creation [park2008long, hu2023incentivizing, yao2024unveiling], may effectively foster more equitable creator ecosystems. And to the extent that unequal earnings distributions reflect a platform’s algorithmic market power [o2024algorithmic], antitrust policy may have a role to play too.66endnote: 6But if algorithmic addiction suspends rational consumer choice then consumer protection laws may be a more appropriate algorithmic governance channel.
Despite previous studies employing Patreon data [regner2021crowdfunding, el2022quantifying, hutson2021youtube], we provide the first systematic analysis of the earnings distribution’s underlying structure and the economic processes or algorithmic mechanisms generating it. We also provide some of the only evidence on content creator earnings distributions [influencermarketinghub2023creator, peres2024creator, Naceva_2025_CreatorEP].
Section 2 presents our data and power law methodology, including the “rich-get-richer” dynamic underpinning it (discussed further in Appendices). Section 3 presents our findings. Section 4 discusses policy implications: how to incentivize greater incorporation of long-tail content [anderson2012long] in recommendation systems, and the impact of market power on this. Section 5 concludes. Robustness and data imputation is discussed in Appendix A.
@@ 2  Data and Method
This section motivates our use of a power law to estimate earnings, and our interpretation of it as reflecting a ‘rich-get-richer’ dynamic, drawing on the preferential attachment framework introduced by [barabasi1999emergence] and its BA-style mixed (copying) variants [mitzenmacher2004brief, easley2010networks, chung2003generalizations]. These models predict that attention follows a power law distribution with density $p(A)\propto A^{-\alpha}$; in a simple mixed copying model, the limiting case of pure preferential attachment ($\gamma=0$) yields a heavy-tailed distribution with $\alpha=2$. Before discussing the model we first detail our data.
@@ 2.1  Data and Descriptive Statistics
We use official Patreon monthly earnings data for content creators purchased directly from the Patreon platform, for the month of March for years 2018, 2021, and 2024 (n = 104,719).
Patreon is a monetization platform for content creators, allowing them to sell additional or primary content to loyal subscribers (paid and unpaid). This is done by linking to their Patreon from their existing social media profiles. Viewers discover these links when they encounter the creator’s content or visit their profile on the third-party social media site, such as Instagram or YouTube. Patreon typically appeals to creators with close follower relationships. Other common link-driven monetization strategies by creators on platforms include external links to Amazon, Spotify, Etsy, or App Stores, cross-promotion of the creator’s other content on the same platform, or using internal platform monetization tools [rieder2023making].
As a result Patreon earnings are highly unlikely to be the sole type of earnings for these content creators, since they can also make money directly from their affiliation with the social media platform, e.g. from YouTube affiliates, ads, and sponsorships, unless their only affiliation is Patreon. But this data is likely to be as unequally distributed – if not more, since it appears to scale exponentially with viewership [influencermarketinghub2023creator]. Unfortunately, there is no detailed public data on this.
Patreon, however, allows creators to keep their earnings private, leading to significant missing data here, especially for larger earners – since those with more paid subscribers tend to disclose their earnings the least, we find. Not imputing missing earnings, therefore, would introduce a significant downward bias on the right tail of the distribution. We impute missing earnings linearly using a basic specification, based on earnings being a function of paid members (subscribers), all members, the type of content created, if they are an NSFW content creator, and the year of the data. We show in an Appendix C that this is the method is robust and has good predictive power, despite its simplicity. We then restrict the dataset to include only content creators with monthly earnings exceeding \$10 USD.
Next, we segment the data by the social media platform(s) the Patreon content creators use to engage their audiences. We identify creators who only use a single social media platform and remove from our dataset all earners with more than one social media affiliation. The dataset spans Patreon creators across major platforms, including YouTube, Instagram, Twitter, Twitch, Facebook, and Patreon only. Table 1 provides summary statistics for earnings by platform, detailing the number of observations, mean earnings, standard deviation (SD), and key percentiles (minimum, 25th, 50th, 75th, and maximum).
Table 1: Summary Statistics: Patreon monthly earnings by social media platform
@@
Facebook 8,458 \$149 \$47 \$631 \$10 \$30 \$47 \$99 \$35,261
Instagram 35,879 \$226 \$59 \$1,076 \$10 \$33 \$59 \$142 \$60,391
Patreon 83,950 \$278 \$57 \$1,667 \$10 \$31 \$57 \$154 \$211,321
Twitch 1,793 \$198 \$46 \$1,174 \$10 \$29 \$46 \$105 \$31,436
Twitter 47,564 \$373 \$72 \$2,159 \$10 \$34 \$72 \$215 \$169,106
Youtube 23,435 \$250 \$47 \$1,490 \$10 \$28 \$47 \$119 \$87,802
@@
The data show substantial variation in earnings across platforms. Twitter and Patreon have the highest mean, median, maximum and interquartile earnings in our sample, while Facebook and Twitch sit at the lower end. Twitter and Patreon have the highest mean, median, maximum, and interquartile range of earnings in our sample, while Facebook and Twitch sit at the lower end. At least part of this pattern, especially the very high maxima, is likely driven by the much larger number of creators we observe on Twitter and Patreon, rather than by systematically different underlying inequality across platforms.77endnote: 7Raw maxima and interquartile ranges can both be misleading indicators of inequality when outcomes have very heavy upper tails, as is the case for creator earnings. Intuitively, a platform with more creators has many more “draws from the lottery” of extreme success. Even if two platforms had identical underlying earnings distributions, the one with more creators would tend to exhibit a higher observed maximum simply because it is more likely to contain an extremely successful creator. In the following sections we show that creator earnings are well described by a power-law (Pareto-type) distribution. In such distributions, the expected maximum in a sample of size $N$ grows quickly with $N$ (roughly like $N^{1/(\alpha-1)}$ in a Pareto model), so larger platforms naturally produce more extreme top observations even when the underlying tail parameter $\alpha$, our measure of inequality, is the same. At the same time, the interquartile range focuses on the middle of the distribution and ignores the very top tail, so platforms with very different degrees of top-end concentration can have similar interquartile ranges. This is why, for characterizing inequality, we focus on the estimated exponent $\alpha$ rather than on raw maxima, interquartile ranges, or other platform-specific summary statistics. We note later that this proxy for ecosystem health or strength (median, mean, and interquartile earnings), correlates strongly with the estimated power law for each platform. The standard deviation of earnings is highest for Twitter and lowest for Facebook. Facebook is the only platform to have fewer exclusive Patreon content creators in our data in 2024 than it had in 2021 (Table 2 in Appendix).
@@ 2.2  Power Law Model
Power law distributions, $P(X)\sim\frac{1}{x^{\alpha}}$, provide a useful framework for characterizing earnings online, offering insights into system scaling, earnings inequality, and underlying mechanisms. These distributions describe systems where small occurrences are extremely common while large occurrences are extremely rare but have significant impact. A lower power law exponent implies higher inequality, indicating greater probability of extreme tail incomes characteristic of long-tail distributions such as wealth [pareto1896cours, atkinson1970distribution] or internet traffic, where a few websites capture most visits [adamic2002zipf, huberman1998internet].
Much evidence shows that various aspects of society and the economy follow power laws [gabaix2009power], with the vast majority of outcomes small yet a few extraordinarily large, commanding substantial shares of the total. Online markets often exhibit similar extreme disparities, with a small number of top content creators or e-commerce sellers capturing disproportionate revenue shares [fortunato2006topical, clauset2009power].
The hypothesis we estimate is that “rich-get-richer” dynamics, characterized by preferential attachment models [barabasi1999emergence], govern content creator earnings. As creators gain visibility, they attract further attention through algorithmic amplification [zappin2022youtube, hua2022characterizing], creating a self-reinforcing feedback loop that concentrates income among a small group of highly popular creators – a “winner takes all” or viral effect [rosen1981economics, rieder2018ranking]. This compounding of initial advantage, also called a “superstar effect” in economics [gabaix2009power, autor2020fall], means that participants who accumulate more engagement become even more likely to attract future interactions. Patreon earnings data provide concrete, measurable outcomes of these dynamics given their reliance on third-party platform algorithms for visibility.
A similar dynamic of persistence governs firm dynamics in the broader economy [geroski2001modelling, geroski2002learning, bottazzi2006explaining]; however, online, network and algorithmic effects can supercharge this [easley2010networks].88endnote: 8The low marginal cost to distributing and selling content online adds to the power law, which arises from the feedback introduced by correlated decisions across a population [easley2010networks].99endnote: 9Some econophysics models distinguish a two-class structure in income distributions, where working class income follows an exponential distribution (reflecting additive processes) and rentier or capitalist class income follows a power law (reflecting multiplicative processes). See [ludwig2022physics, druagulescu2001exponential].
Detecting power law distributions in earnings data provides valuable insights into the health and fairness of online ecosystems. Highly skewed distributions, with large numbers earning very little, may signal systemic challenges such as algorithmic bias, barriers to entry, or lack of upward mobility [anderson2022patreon, needleman2023influencers]. Conversely, distributions less dominated by top earners suggest more inclusive and competitive environments with broader opportunities for success. Understanding these patterns can inform platform governance, guide policymakers, and help researchers propose fairer algorithms and revenue-sharing mechanisms, ultimately leading to more equitable digital economies [eckhardt2023creator, gillespie2018custodians].